# Are We Asking the Right Questions?: Designing for Community Stakeholders’ Interactions with AI in Policing

### 08 Feb 2024 v1

@@ 2.3. Community Members’ Participation in AI Deliberation
Academics and journalists have gathered substantial evidence highlighting the outcome-oriented harms resulting from algorithmic systems and have specifically drawn attention to disparate harms inflicted on impacted communities (raji2022fallacy, ; shelby2022identifying, ; johnson_2023, ; Villasenor_2021, ; garance_2022, ; garza_2022, ). Consequently, the design process of AI systems is often criticized for not involving community members in the process even though they are the ones who face the consequences of algorithmic decisions. As a result, researchers have argued for expanding participation in the AI deliberation and design process to incorporate a variety of stakeholders and elicit their feedback at different stages of the process (delgado2023participatory, ). SIGCHI researchers have conducted studies that elicit community members’ perceptions of fairness (van2021effect, ; van2019crowdsourcing, ; srivastava2019mathematical, ) as well as designed frameworks to elicit stakeholders’ nuanced notions of fairness regarding algorithmic systems deployed in real-world contexts (cheng2021soliciting, ). Similarly, researchers have engaged in co-design processes with workers to develop AI interventions that are centered in their practice and foster worker well-being (holstein2019co, ; zhang2022algorithmic, ; spektor2023designing, ).
@@
Regarding algorithmic systems deployed in the public sector, researchers have closely worked with civil servants whose labor is impacted by these systems and elicited their feedback to inform the ethical and human-centered design of AI systems (brown2019toward, ; cheng2022child, ; saxena2020conducting, ; ammitzboll2021street, ). Recently, SIGCHI researchers have contributed to this literature by working directly with impacted community members to understand their perspectives and needs regarding the use of AI systems in public agencies and NGOs. For instance, Lee et al. (lee2019webuildai, ) developed a participatory framework that facilitates ML model development with a variety of stakeholders and improves distributive outcomes and procedural fairness. Brown et al. (brown2019toward, ) conducted participatory workshops with community members to understand their perspectives on the use of a predictive risk model (PRMs) deployed within the child welfare system. Stapleton et al. (stapleton2022imagining, ) build upon this work by conducting design workshops with parents investigated by the child welfare system and social workers to question whether such models should be developed, to begin with, and found that both these stakeholders had significant concerns regarding PRMs and instead wanted to address the problems motivating the use of such systems. Similarly, Kuo and Shen et al. (kuo2023understanding, ) engaged stakeholders, including service providers, county workers, and unhoused individuals, in AI lifecycle comicboarding sessions for homeless services. Their findings indicate that stakeholders can actively contribute to the early stages of the AI deliberation process, influencing problem formulation, dataset selection, and modeling choices. This study further builds upon this body of research by introducing an intuitive AI application that acts as a boundary object that the participants can interact with, helps facilitate context-specific and nuanced conversations with them about policing practices in their city, the use of AI systems by police departments, and elicit their feedback regarding the ethical design and use of such systems.
Figure 1. A simplified sequence of events in the criminal justice system.
@@ 3. Methods
We conducted a mixed methods study with participants from Milwaukee, Wisconsin, a mid-sized, midwestern city in the United States from May to August 2019. A total of 60 community stakeholders participated in this study and helped us examine how people from different educational and professional backgrounds interact with crime mapping algorithms. Based on prior literature about common crime mapping algorithms used in police departments (chainey_utility_2008, ), we developed an interactive application that displays a heatmap output of a standard Kernel Density Estimation (KDE) algorithm given crime data and particular user inputs. Our study consisted of 2 parts - 1) through an in-lab activity, we examined participants’ interactions with the crime mapping algorithm where they chose different parameters for the KDE algorithm, interpreted the results, and identified hotspots on the map. After the lab activity, participants completed the NASA-TLX survey which helped us assess their self-reported mental workload as a result of the activity, 2) we conducted semi-structured interviews with the participants to gather their perspectives and needs regarding the ethical design and use of such systems by police departments. This study was approved by the Institutional Review Board (IRB) of our university in the US.
@@ 3.1. Application Design
Crime data used in our application was taken from the Wisconsin Incident-Based Reporting System (WIBRS), a publicly accessible database organized by the City of Milwaukee and Milwaukee Police Department (wibrs, ). Data had to be accessed by time and was collected starting with 2017 (the last full year before this project began) and then worked backward. We were able to curate a dataset of crimes from 2012 to 2017. We chose to focus on 3 distinct quality-of-life crimes - robbery, larceny, and motor theft because they are most commonly reported and tend to have the most effect on an average person’s life (james2008crime, ). Moreover, city residents would also be familiar with or might have encountered them. In contrast, rape, murder, terrorism, etc. receive significant media coverage but are comparatively rare events and also disproportionately reported between rural and urban settings (livingston1997police, ). Street addresses provided in the dataset were then cleaned using a manually assembled regex library. Then they were geocoded using Google’s geocoding service, and organized into individual months.
To facilitate our study, we built an online, interactive application (Figure 2) that displays the heatmap output of a Bivariate Kernel Density Estimation (KDE) based on specified data and KDE parameters. We chose to present KDE outputs for two reasons: first, we knew from prior knowledge of common practices by crime labs that KDEs are one of, if not the most commonly used crime mapping algorithm (haque2019exploring, ), and second, we wanted to examine participants’ interactions with and interpretation of the KDE heatmap. Each parameter dropdown list had a question mark tooltip to answer some general questions about parameters. The data parameters allow users to choose one type of crime from a specific month from 2012 to 2017. To the extent possible, we chose to develop the design features of our application that closely resemble popular crime mapping platforms used in police departments. Moreover, such crime mapping platforms are usually not accessible to researchers so we decided to create an in-house application that closely mimics a real-world algorithmic crime mapping application. KDE parameters and options are outlined in Table 1.
@@
Kernel Gaussian Tophat, Epanechnikov, Linear
Bandwidth 0.005 0.0025, 0.0075, 0.01
Distance Metric Euclidean Manhattan, Chebyshev
Table 1. List of Interactive KDE Parameters and Options for Participants
@@ 3.2. How does KDE Map Crime?
KDEs attempt (backurs2019space, ) to find and represent the underlying probability distribution function (PDF) that a set of data was taken from. By providing a PDF, the heatmap allows users to predict high- and low-probability areas for future events. KDEs accomplish this by smoothing each discrete data point into a two-dimensional probability distribution function with the original point at the mean and aggregating the PDFs into a singular heatmap for the entire area. The shape of the distribution that data points are smoothed into is defined by the kernel parameter, also called the interpolation method. The bandwidth parameter controls the width of each distribution. In statistics, this would be analogous to the variance of a symmetric distribution. For example, a Gaussian kernel with a bandwidth of one will result in the standard normal distribution, though generally a much smaller bandwidth is preferred in order to produce meaningful results. Higher values for bandwidth result in much smoother outputs, which would lose power but reduce bias from overfitting. The distance metric parameter controls how the distance between points is measured. The most commonly used distance metric is Euclidean, or merely a straight line between two points on a plane, which is the default setting for the distance metric. In our application, after submitting an initial set of parameters, users are shown the calculated heatmap with several interactive features, as well as a bar chart displaying the number of events in either each police district or each aldermanic district. Hovering over data points in the heatmap shows a pop-up with information about the date, district, and location of the event.
@@ 3.3. Participants and Data Collection
We decided to recruit participants local to this midwestern city as most empirical studies on algorithms in criminal justice either deal with historical data about people suspected or convicted of crime or crowd workers (e.g. Amazon Mechanical Turk, Crowdflower, etc.) (barabas2019technical, ; rubya2021comparing, ; sen2015turkers, ; lee2021included, ). Local residents are able to situate the results and interpret them in the context of their lived experiences allowing us to better understand their perspectives and needs.
Recruitment was undertaken in the form of advertisement flyers posted around the city in coffee shops, community centers, and public libraries as well as various online forums and groups. As much as possible, we tried to disperse our recruitment efforts through all parts of the city including high- and low-income areas as well as those parts of the city with historically higher or lower rates of crime. To recruit LEA participants, we made special efforts to reach out to various police departments in the county which included about 8 different police departments. While we were successful in recruiting LEA participants, this was significantly hard because LEAs in these departments are busy, under-resourced, and generally reticent to speak to researchers as there is a history of critical research and policy findings about the practices of police departments (july15july20, ). In the end, we ended up with participants ($n=60$) with varying academic and professional backgrounds. An overview of the participants’ demographics can be found in Table 2. We were able to roughly align gender and education with the 2010 US Census (bureau, ) but were unable to do so for age. Thus, our sample age distribution skews about 20% younger than in the 2010 US Census but is a more relevant and local sample than prior work. Each participant was given a specific participant identification number to allow us to use their data and interview transcript for analysis without revealing personally identifiable information.
@@ 3.4. Participant Grouping
Demographic criteria Participant description (with no. of participants)
Gender Male (23); Female (36); Trans male (1)
Age range (in years) 18-21 years (29); 22-30 years (20); 31-40 years (4); 40+ years (7)
@@
High school diploma or GED (2); Undergraduate (32);
Bachelors (20); Masters (5); Doctorate (1)
@@
Full-time (17); part-time (14); unemployed (4);
self-employed (3); full-time student (21); retired (1)
@@ 3.5. In-Lab Activities
Following this, participants were asked to envision green circles with crosshairs as the optimal patrol areas for one police unit. They were then tasked with specifying the minimum and maximum number of circles required to address crime effectively based on the heatmap. Finally, they placed 5 circles in locations representing the most efficient allocation of resources. An example heatmap can be seen in Figure 2. The session, led by one author and assisted by another, involved these interactive lab activities. Post-session, the authors collaborated to refine observation notes and shared these impressions. Participants then completed the NASA-TLX survey (hart1988development, ) to evaluate mental workload, a standard instrument for assessing the efficacy of a system in HCI (NASA, ). Participants then answered a survey with questions about their prior level of familiarity with algorithmic crime analysis, their understanding of the application, feelings towards law enforcement and government, and general demographics. Finally, participants were interviewed with a set of questions in a protocol based on their familiarity with such algorithmic systems, their views on the legality, ethics, and fairness of data, and their concerns regarding algorithmic crime mapping. The interview sessions were audio-recorded with participants’ consent.
@@ 3.6. Analysis
We calculated the probability distribution of using the default parameters to determine the default behavior among the three groups. Additionally, we used the NASA-TLX survey to assess the mental workload of interacting with the crime map. The TLX involves participants reflecting on the task and making 15 paired comparisons across six dimensions to gauge workload. For instance, they needed to decide whether Performance or Frustration represented the more important contributor to the workload for the specific task they had recently performed. The second step involved participants rating each of the six dimensions on a scale from Low to High or from Good to Poor. The raw score for each of the six items is multiplied by the weight from step 1 to generate the overall workload score per task.
Qualitative interviews were transcribed using the online platform Otter (otter, ). Subsequently, two authors manually reviewed the automated transcriptions to ensure accuracy. Thematic analysis (braun2006using, ) with constant comparison was employed, and two authors conducted iterative coding of textual codes. A third co-author verified the coding process. These low-level codes were combined into higher-level categories by combining all three authors’ suggestions. Finally, these categories were collated into higher-level themes using our knowledge of prior work as well as practical, domain knowledge about criminal justice systems. These themes are presented throughout our results below with appropriate quotes and explanations.
@@ 4. Results
Before engaging in the lab activities, we gave participants an overview of algorithmic crime mapping. We shared with them how the application was expected to work and the general intent behind algorithmic crime mapping (i.e., efficient allocation of resources on the part of the police department). We also showed them how they could produce different maps by changing the parameters. After observing their interactions with the application, we conducted semi-structured interviews to gather their impressions of this technology and learn about their needs. To improve the readability of the results, we first share participants’ perspectives (RQ1), followed by the findings from participants’ interactions with the applications (RQ2), and finally, their needs regarding a community-centered vision of this application (RQ3).
@@ 4.1. Participants’ Perspectives on Algorithmic Crime-Mapping (RQ1)
Most community members and technical participants shared their skepticism about crime mapping where they were primarily concerned that such technologies could lead to over-policing of specific regions, used as a revenue generator by targeting specific areas, and create a feedback loop that leads to over-policing of specific neighborhoods and communities. Several participants questioned the motivation behind crime mapping and wondered if the technology was being used to protect citizens and deter crime or whether it was being used simply to identify a high-crime region and target it to generate revenue. For instance, participant P1 asked -
@@
Here, another participant also brought up the example of speed traps as well as other traffic violations and believed that traffic cameras and other forms of automation were better suited to deter those crimes than increasing the police presence in the neighborhood as well as questioned whether increased police presence led to safer neighborhoods. The participant who lives close to an urban college campus shared -
@@
Some technical participants were also concerned about how the application changes over time and whether focusing on certain crimes draws attention (and resources) away from more serious crimes as well as the long-term impact of increased police presence in some neighborhoods. For instance, T5 shared -
@@
On the other hand, several community members and law enforcement officers found utility in the tool. Community members appreciated that such applications could provide a real-time snapshot of crime prevalence in different neighborhoods and help them make important decisions such as purchasing a house and finding the right school district for their children. Participant, P8, went so far as to say that they would be willing to freely share their data to create such a public resource.
@@
Law enforcement officers also shared that they found utility in such a tool because it helped them make real-time data-driven decisions based on crime statistics in certain areas and decide how many patrol cars they needed on the streets. For instance, L8 shared -
@@
In sum, we learned that several community members found utility in algorithmic crime mapping but were primarily concerned about the motivations driving the use of such tools. They shared that instead of protecting people and deterring crime, the tool might become a way for police officers to “find crime” and set up people as a means for writing more tickets. Here, it is important to note that the distrust in policing practices underscored the majority of our conversations with community members who often referred to prominent cases of police brutality to highlight police wrongdoing. As we highlight in Section 4.3, several community members brought up examples where the police reports did not match the video evidence shared on social media by bystanders of the same incidents.

### 19 Mar 2024 v2

@@ 2.3. Community Members’ Participation in the AI Deliberation Process
Academics and journalists have gathered substantial evidence highlighting the outcome-oriented harms resulting from algorithmic systems and have specifically drawn attention to disparate harms inflicted on impacted communities (Raji et al., 2022; Shelby et al., 2022; Johnson, 2023; Villasenor and Foggo, 2021; Ho and Burke, 2022; De La Garza, 2020). Consequently, the design process of AI systems is often criticized for not involving community members in the process even though they are the ones who face the consequences of algorithmic decisions. As a result, researchers have argued for expanding participation in the AI deliberation and design process to incorporate a variety of stakeholders and elicit their feedback at different stages of the process (Delgado et al., 2023). SIGCHI researchers have conducted studies that elicit community members’ perceptions of fairness (Van Berkel et al., 2021, 2019; Srivastava et al., 2019) as well as designed frameworks to elicit stakeholders’ nuanced notions of fairness regarding algorithmic systems deployed in real-world contexts (Cheng et al., 2021). Similarly, researchers have engaged in co-design processes with workers to develop AI interventions that are centered in their practice and foster worker well-being (Holstein et al., 2019; Zhang et al., 2022; Spektor et al., 2023).
@@
Regarding algorithmic systems deployed in the public sector, researchers have closely worked with civil servants whose labor is impacted by these systems and elicited their feedback to inform the ethical and human-centered design of AI systems (Brown et al., 2019; Cheng et al., 2022; Saxena and Guha, 2020; Ammitzbøll Flügge et al., 2021). Recently, SIGCHI researchers have contributed to this literature by working directly with impacted community members to understand their perspectives and needs regarding the use of AI systems in public agencies and NGOs. For instance, Lee et al. (Lee et al., 2019) developed a participatory framework that facilitates ML model development with a variety of stakeholders and improves distributive outcomes and procedural fairness. Brown et al. (Brown et al., 2019) conducted participatory workshops with community members to understand their perspectives on the use of a predictive risk model (PRMs) deployed within the child welfare system. Stapleton et al. (Stapleton et al., 2022a) build upon this work by conducting design workshops with parents investigated by the child welfare system and social workers to question whether such models should be developed, to begin with, and found that both these stakeholders had significant concerns regarding PRMs and instead wanted to address the problems motivating the use of such systems. Similarly, Kuo and Shen et al. (Kuo et al., 2023) engaged stakeholders, including service providers, county workers, and unhoused individuals, in AI lifecycle comicboarding sessions for homeless services. Their findings indicate that stakeholders can actively contribute to the early stages of the AI deliberation process, influencing problem formulation, dataset selection, and modeling choices. This study further builds upon this body of research by introducing an intuitive AI application that acts as a boundary object that the participants can interact with, helps facilitate context-specific and nuanced conversations with them about policing practices in their city, the use of AI systems by police departments, and elicit their feedback regarding the ethical design and use of such systems.
@@ 3. Methods
We conducted a mixed methods study with participants from Milwaukee, Wisconsin, a mid-sized, midwestern city in the United States from May to August 2019. A total of 60 community stakeholders participated in this study and helped us examine how people from different educational and professional backgrounds interact with crime mapping algorithms. Based on prior literature about common crime mapping algorithms used in police departments (Chainey et al., 2008), we developed an interactive application that displays a heatmap output of a standard Kernel Density Estimation (KDE) algorithm given crime data and particular user inputs. Our study consisted of 2 parts - 1) through an in-lab activity, we examined participants’ interactions with the crime mapping algorithm where they chose different parameters for the KDE algorithm, interpreted the results, and identified hotspots on the map. After the lab activity, participants completed the NASA-TLX survey which helped us assess their self-reported mental workload as a result of the activity, 2) we conducted semi-structured interviews with the participants to gather their perspectives and needs regarding the ethical design and use of such systems by police departments. This study was approved by the Institutional Review Board (IRB) of our university in the United States.
@@ 3.1. Application Design
Crime data used in our application was taken from the Wisconsin Incident-Based Reporting System (WIBRS), a publicly accessible database organized by the City of Milwaukee and Milwaukee Police Department (Author, 2019b). Data had to be accessed by time and was collected starting with 2017 (the last full year before this project began) and then worked backward. We were able to curate a dataset of crimes from 2012 to 2017. We chose to focus on 3 distinct quality-of-life crimes - robbery, larceny, and motor theft because they are most commonly reported and tend to have the most effect on an average person’s life (James, 2008). Moreover, city residents would also be familiar with or might have encountered them. In contrast, rape, murder, terrorism, etc. receive significant media coverage but are comparatively rare events and also disproportionately reported between rural and urban settings (Livingston, 1997). Street addresses provided in the dataset were then cleaned using a manually assembled regex library. Then they were geocoded using Google’s geocoding service, and organized into individual months.
To facilitate our study, we built an online, interactive application (Figure 1) that displays the heatmap output of a Bivariate Kernel Density Estimation (KDE) based on specified data and KDE parameters. We chose to present KDE outputs for two reasons: first, we knew from prior knowledge of common practices by crime labs that KDEs are one of, if not the most commonly used crime mapping algorithm (Haque et al., 2019), and second, we wanted to examine participants’ interactions with and interpretation of the KDE heatmap. Each parameter dropdown list had a question mark tooltip to answer some general questions about parameters. The data parameters allow users to choose one type of crime from a specific month from 2012 to 2017. To the extent possible, we chose to develop the design features of our application that closely resemble popular crime mapping platforms used in police departments. Moreover, such crime mapping platforms are usually not accessible to researchers so we decided to create an in-house application that closely mimics a real-world algorithmic crime mapping application. KDE parameters and options are outlined in Table 1.
@@
Kernel
Gaussian
Tophat, Epanechnikov, Linear
Bandwidth
0.005
0.0025, 0.0075, 0.01
Distance Metric
Euclidean
Manhattan, Chebyshev
Table 1. List of Interactive KDE Parameters and Options
@@ 3.2. How does KDE Map Crime?
KDEs attempt (Backurs et al., 2019) to find and represent the underlying probability distribution function (PDF) that a set of data was taken from. By providing a PDF, the heatmap allows users to predict high- and low-probability areas for future events. KDEs accomplish this by smoothing each discrete data point into a two-dimensional probability distribution function with the original point at the mean and aggregating the PDFs into a singular heatmap for the entire area. The shape of the distribution that data points are smoothed into is defined by the kernel parameter, also called the interpolation method. The bandwidth parameter controls the width of each distribution. In statistics, this would be analogous to the variance of a symmetric distribution. For example, a Gaussian kernel with a bandwidth of one will result in the standard normal distribution, though generally a much smaller bandwidth is preferred in order to produce meaningful results. Higher values for bandwidth result in much smoother outputs, which would lose power but reduce bias from overfitting. The distance metric parameter controls how the distance between points is measured. The most commonly used distance metric is Euclidean, or merely a straight line between two points on a plane, which is the default setting for the distance metric. In our application, after submitting an initial set of parameters, users are shown the calculated heatmap with several interactive features, as well as a bar chart displaying the number of events in either each police district or each aldermanic district. Hovering over data points in the heatmap shows a pop-up with information about the date, district, and location of the event.
@@ 3.3. Participants and Data Collection
We decided to recruit participants local to this midwestern city as most empirical studies on algorithms in criminal justice either deal with historical data about people suspected or convicted of crime or crowd workers (e.g. Amazon Mechanical Turk, Crowdflower, etc.) (Barabas et al., 2019; Rubya et al., 2021; Sen et al., 2015; Lee and Rich, 2021). Local residents are able to situate the results and interpret them in the context of their lived experiences allowing us to better understand their perspectives and needs.
Recruitment was undertaken in the form of advertisement flyers posted around the city in coffee shops, community centers, and public libraries as well as various online forums and groups. As much as possible, we tried to disperse our recruitment efforts through all parts of the city including high- and low-income areas as well as those parts of the city with historically higher or lower rates of crime. To recruit LEA participants, we made special efforts to reach out to various police departments in the county which included about 8 different police departments. While we were successful in recruiting LEA participants, this was significantly hard because LEAs in these departments are busy, under-resourced, and generally reticent to speak to researchers as there is a history of critical research and policy findings about the practices of police departments (Author, 2020). In the end, we ended up with participants ($n=60$) with varying academic and professional backgrounds. An overview of the participants’ demographics can be found in Table 2. We were able to roughly align gender and education with the 2010 US Census (Bureau, 2019) but were unable to do so for age. Thus, our sample age distribution skews about 20% younger than in the 2010 US Census but is a more relevant and local sample than prior work. Each participant was given a specific participant identification number to allow us to use their data and interview transcript for analysis without revealing personally identifiable information.
@@ 3.4. Participant Grouping
Demographic Criteria
Participant Description (with number of participants)
Gender
Male (23); Female (36); Trans male (1)
Age range
18-21 years (29); 22-30 years (20); 31-40 years (4); 40+ years (7)
@@
High school diploma or GED (2); Undergraduate (32); Bachelors (20); Masters (5); Doctorate (1)
@@
Full-time (17); part-time (14); unemployed (4); self-employed (3); full-time student (21); retired (1)
@@ 3.5. In-Lab Activities
Following this, participants were asked to envision green circles with crosshairs as the optimal patrol areas for one police unit. They were then tasked with specifying the minimum and maximum number of circles required to address crime effectively based on the heatmap. Finally, they placed 5 circles in locations representing the most efficient allocation of resources. An example heatmap can be seen in Figure 1. The session, led by one author and assisted by another, involved these interactive lab activities. Post-session, the authors collaborated to refine observation notes and shared these impressions. Participants then completed the NASA-TLX survey (Hart and Staveland, 1988) to evaluate mental workload, a standard instrument for assessing the efficacy of a system in HCI (Prabaswari et al., 2019). Participants then answered a survey with questions about their prior level of familiarity with algorithmic crime analysis, their understanding of the application, feelings towards law enforcement and government, and general demographics. Finally, participants were interviewed with a set of questions in a protocol based on their familiarity with such algorithmic systems, their views on the legality, ethics, and fairness of data, and their concerns regarding algorithmic crime mapping. The interview sessions were audio-recorded with participants’ consent.
@@ 3.6. Analysis
We calculated the probability distribution of using the default parameters to determine the default behavior among the three groups. Additionally, we used the NASA-TLX survey to assess the mental workload of interacting with the crime map. The TLX involves participants reflecting on the task and making 15 paired comparisons across six dimensions to gauge workload. For instance, they needed to decide whether Performance or Frustration represented the more important contributor to the workload for the specific task they had recently performed. The second step involved participants rating each of the six dimensions on a Likert-type scale. The raw score for each of the six items is multiplied by the weight from step 1 to generate the overall workload score per task.
Qualitative interviews were transcribed using the online platform Otter (otter, [n. d.]). Subsequently, two authors manually reviewed the automated transcriptions to ensure accuracy. Thematic analysis (Braun and Clarke, 2006) with constant comparison was employed, and two authors conducted iterative coding of textual codes. A third co-author verified the coding process. These low-level codes were combined into categories, and finally, these categories were collated into higher-level themes. These themes are presented throughout our results below with appropriate quotes and explanations.
@@ 4. Results
Before engaging in the lab activities, we provided the participants with an overview of algorithmic crime mapping. We shared with them how the application was expected to work and the general intent behind algorithmic crime mapping (i.e., efficient allocation of resources on the part of the police department). We also showed them how they could produce different maps by changing the parameters. After observing their interactions with the application, we conducted semi-structured interviews to gather their impressions of this technology and learn about their needs. To improve the readability of the results, we first share participants’ perspectives (RQ1), followed by the findings from participants’ interactions with the applications (RQ2), and finally, their needs regarding a community-centered vision of this application (RQ3).
@@ 4.1. Participants’ Perspectives on Algorithmic Crime-Mapping (RQ1)
Most community members and technical participants shared their skepticism about crime mapping where they were primarily concerned that such technologies could lead to over-policing of specific regions, used as a revenue generator by targeting specific areas, and create a feedback loop that leads to over-policing of specific neighborhoods and communities. Several participants questioned the motivation behind crime mapping and wondered if the technology was being used to protect citizens and deter crime or whether it was being used simply to identify a high-crime region and target it to generate revenue. For instance, participant C1 asked:
@@
Here, another participant also brought up the example of speed traps as well as other traffic violations and believed that traffic cameras and other forms of automation were better suited to deter those crimes than increasing the police presence in the neighborhood as well as questioned whether increased police presence led to safer neighborhoods. The participant who lives close to an urban college campus shared:
@@
Some technical participants were also concerned about how the application changes over time and whether focusing on certain crimes draws attention (and resources) away from more serious crimes as well as the long-term impact of increased police presence in some neighborhoods. For instance, T5 shared:
@@
On the other hand, several community members and law enforcement officers found utility in the tool. Community members appreciated that such applications could provide a real-time snapshot of crime prevalence in different neighborhoods and help them make important decisions such as purchasing a house and finding the right school district for their children. Participant, C6, went so far as to say that they would be willing to freely share their data to create such a public resource.
@@
Law enforcement officers also shared that they found utility in such a tool because it helped them make real-time data-driven decisions based on crime statistics in certain areas and decide how many patrol cars they needed on the streets. For instance, L8 shared:
@@
In sum, we learned that several community members found utility in algorithmic crime mapping but were primarily concerned about the motivations driving the use of such tools. They shared that instead of protecting people and deterring crime, the tool might become a way for police officers to “find crime” and set up people as a means for writing more tickets. Here, it is important to note that the distrust in policing practices underscored the majority of our conversations with community members who often referred to prominent cases of police brutality to highlight police wrongdoing. As we highlight in Section 4.3, several community members brought up examples where the police reports did not match the video evidence shared on social media by bystanders.